Скачать презентацию Weekly Review of Operation Recycler departmental meeting May Скачать презентацию Weekly Review of Operation Recycler departmental meeting May

552fa245b1f8c0832ededda54e88f0cf.ppt

  • Количество слайдов: 14

Weekly Review of Operation Recycler departmental meeting May 23 rd, 2007 Weekly Review of Operation Recycler departmental meeting May 23 rd, 2007

Agenda § Machine status 2 Agenda § Machine status 2

Status Summary – Transfers from Accumulator § Pbars to the Recycler Ø 43 Shot Status Summary – Transfers from Accumulator § Pbars to the Recycler Ø 43 Shot setups • 1829 E 10 Pbars extracted from the Accumulator (in MI) – Two months ago, same number of shot setups → 2200 e 10 ! – Assuming ‘Historical’ MI to RR efficiencies (i. e. 95%): 1738 e 10 – Last week: 1770 e 10 » Average efficiencies Historical – ACC to MI 95% Last week This week XX% 93% Note: Two points lower for 3 -day average following Accumulator lattice change and one point higher after that » ‘Storage’ XX% 82% 85% (in/out measured in MI) Did not take into account ‘storage cycle’ where we lost 150 e 10 • for that cycle: Storage efficiency was 41% 3

Pbar Monthly Totals 4 Pbar Monthly Totals 4

Status Summary – Shots to Te. V § Pbars to the Tev Ø 6 Status Summary – Shots to Te. V § Pbars to the Tev Ø 6 shot setups: 1445 E 10 transferred (measured in MI) • 241 e 10 Average Intensity per shot – Last week: 245 e 10; 2 months ago: 300 e 10 Ø Average emittances – Trans. Schottky (in RR)* – Trans. FW (in RR) * [p mm mrad] – Longitudinal (in RR) * – Long. (bunch in MI) [e. V s] • Average mining efficiency : Last week 5. 0 1. 8 56 XX This week 4. 9 2. 1 59 1. 0 Note: I am now taking the average of all the bunches XX% 96% * Before mining 5

Operational Highlights and Machine Maintenance § Recyler Ø LLRF module replaced (VXI card) • Operational Highlights and Machine Maintenance § Recyler Ø LLRF module replaced (VXI card) • Fixed the slope we were seeing a week or so ago Ø Adjusted injection phase by 1 bucket (this morning) Ø Indications that the R 32 line is getting close to a ‘natural’ vacuum burst (12 -15 weeks cycle Argon instability) • Tried to induce it twice by closing/opening valves, turning ion pumps on/off and degassing a couple gauges – No luck § Ecool Ø Increased the time needed for the MCC channel to be low before inducing a trip • This is to avoid MCC trips due to electrical noise Ø Adjusted MI-31 temperature after thermostat batteries were replaced • i. e. was effectively cooling Ø Conditioned the Pelletron (sections 1 thru 4) • Running well since 6

Issues - Recycler § Lost 30 e 10 during transfers Ø Beam was over Issues - Recycler § Lost 30 e 10 during transfers Ø Beam was over squeezed • DC beam kicked out by injection kicker § Lost 120 e 10 when QCLP tripped Ø Gain was being changed • QCL ramp did something funny when original setting was restored and tripped § The stochastic cooling system switches itself to ‘gated mode’ regularly (~3 times per 24 hours) Ø This is not new… anything we can do about it ? 7

Issues - Ecool § ECOOL VME hung up AGAIN !! (Sunday-Tuesday combo) Ø Both Issues - Ecool § ECOOL VME hung up AGAIN !! (Sunday-Tuesday combo) Ø Both times, it required to go into MI-31 pit to power cycle the crate Ø On Sunday, the VME could not reboot at all (even locally) because of a wrong file being loaded • Leftover from changes made the previous Tuesday to monitor/diagnose the issue 8

Studies highlights § Compound bucket during injection w/ gated stochastic cooling Ø Encouraging (more Studies highlights § Compound bucket during injection w/ gated stochastic cooling Ø Encouraging (more data analysis required) ? May be better lifetime too Compound bucket Normal bucket 9

Plan/Upcoming studies § Near term Ø Measure and calibrate cavity fan back voltage (Martin, Plan/Upcoming studies § Near term Ø Measure and calibrate cavity fan back voltage (Martin, Paul, Chandra) Ø Cooling efficiency off-axis (Lionel, Sasha) Ø Momentum diffusion study (Dan) • Need 100 -200 e 10 Ø New optics implementation (Many…) • Wait for latest data analysis ? (Meiqin) Ø Compound bucket with gated stochastic cooling (Chandra, Martin, Sasha, Dan) • With large number of pbars Ø RR dampers update (Jim Crisp, Nathan Eddy) • Still not implemented Ø Adjust compensation loop to new $2 A ramp (Paul Lebrun) • Sometime next week 10

Plan/Upcoming studies (II) Ø Fire TSPs in 800 region (Ron K. , Terry A. Plan/Upcoming studies (II) Ø Fire TSPs in 800 region (Ron K. , Terry A. ) • Requires low number of pbars Ø Fire TSPs under the Pelletron (Ron K. , Terry A. ) • Requires access Ø Replace 5 V PS for stepper motor in MI 31 (Al Legan) • Requires access Ø Test VDPA 1 spare (Julius Lentz) • Requires low number of pbars 11

Long(er) term studies list § Ecool related (Lionel and/or Sasha and/or Arden) Ø Drag Long(er) term studies list § Ecool related (Lionel and/or Sasha and/or Arden) Ø Drag rate measurements Ø Optimize focusing at 190 G Ø Increase beam size in cooling section • First pass not very successful Ø YAG measurements • Beginning of summer shutdown § ‘Old’ items not marked as completed in studies list Ø Ø Rapid pbar transfer scheme (Chandra, Sasha, Martin) Longitudinal emittance growth during cogging (Chandra) Mining studies: mini-barrier buckets (Chandra) Electron cooling in linear bucket (Lionel, Sasha) § Other Ø Standard stochastic cooling rate measurements (Sasha, Lionel, Dan, Cons) Ø Recycler impedance measurement (Martin) 12

Long(er) term studies list (II) § In proton mode Ø R 22 -line dispersion Long(er) term studies list (II) § In proton mode Ø R 22 -line dispersion function measurement (Ming-jen) Ø Mapping LAM 222 field (Ming-jen) Ø Characterize Recycler abort (Dave Johnson) • For Nova project Ø IPM data in boomerang mode (Cons) 13

List of work in progress/scheduled to be done § Things I was aware of… List of work in progress/scheduled to be done § Things I was aware of… Ø Ecool related • MI loss blocking program (Arden) – Done • Pelletron energy stability improvements (Greg, Kermit) – In progress • Automated energy check/drag rate measurement (Arden) • ACCELNET upgrade (Arden) – Computer in house Ø Recycler related • Check chromaticity (Cons, Stan) – Current lattice • Improve longitudinal emittance measurements (Meiqin, Martin, Chandra) • Try to improve leveling at 2. 5 MHz (Martin, Nathan) • Understand RF output when buckets are dropped (Martin) • New DCCT check out (Ron Kellet, Jim Crisp) 14